# Identify As A Human Does: A Pathfinder of Next-Generation Anti-Cheat Framework for First-Person Shooter Games

### v1

@@ 6.1 Incomparability to Prior Works
Table 2: Ablation study results on each of the subsystems.
CT1 System Validation Test
TP TN FP FN Accuracy Recall NPV AUC-ROC OEI2 TP TN FP FN Accuracy Recall NPV AUC-ROC OEI2
Aimbot RevPov 458 4,796 2,901 406 0.614 0.530 0.922 0. 577 1.246 481 5,088 3,075 403 0.616 0.544 0.927 0.584 1.283
RevStats 453 7,290 407 411 0.904 0.524 0.947 0.736 4.941 566 6,797 1,366 318 0.814 0.640 0.955 0.736 2.864
ExSPC 553 6,187 1,510 311 0.787 0.640 0.952 0.779 2.529 558 6,616 1,547 326 0.793 0.631 0.953 0.786 2.586
Hawk 614 5,803 1,894 250 0.750 0.711 0.959 0.732 2.326 642 5,837 2,326 242 0.716 0.726 0.960 0.721 2.126
Wallhack RevPov 573 6,081 2,664 394 0.685 0.593 0.939 0.644 1.670 627 6,364 2,885 398 0.680 0,612 0.941 0.650 1.684
RevStats 604 8,241 504 363 0.911 0.625 0.958 0.783 4.143 855 6,988 2,261 170 0.763 0.834 0.976 0.795 1.946
ExSPC 760 7,258 1,487 207 0.826 0.786 0.972 0.885 3.303 501 8,484 765 524 0.875 0.489 0.942 0.840 3.736
Hawk 798 7,033 1,712 169 0.806 0.825 0.977 0.815 3.118 869 6,902 2,347 156 0.756 0.848 0.978 0.797 2.649
1CT: Cheat type; 2OEI: Operational Efficiency Index.
@@
Table 3: Comparison result on the different model selection of RevStats.
CT1 Model Validation Test
TP TN FP FN Accuracy Recall NPV AUC-ROC OEI2 TP TN FP FN Accuracy Recall NPV AUC-ROC OEI2
Aimbot LogisticRegression 570 6,204 1,493 294 0.791 0.660 0.955 0.733 2.614 713 4,897 3,266 171 0.620 0.807 0.966 0.703 1.772
Naive Bayes 467 6,740 957 397 0.842 0.874 0.944 0.708 3.069 625 5,928 2,235 259 0.724 0.707 0.958 0.717 2.143
RandomForest 415 7,694 3 449 0.947 0.480 0.945 0.740 4.313 427 7,634 529 457 0.891 0.483 0.944 0.709 4.313
SVM 309 7,615 82 555 0.926 0.358 0.932 0.673 7.299 405 7,665 498 479 0.892 0.458 0.941 0.699 4.320
MLP 858 7,697 0 6 0.999 0.993 0.999 0.997 9.901 508 7,325 838 376 0.866 0.575 0.951 0.736 3.674
GaussianNB 467 6,740 957 397 0.842 0.541 0.944 0.708 3.069 625 5,928 2,235 259 0.724 0.707 0.958 0.717 2.143
QDA 350 7,181 516 514 0.880 0.405 0.933 0.669 3.737 534 6,770 1,393 350 0.807 0.604 0.951 0.717 2.697
RevStats 453 7,290 407 411 0.904 0.524 0.947 0.736 4.941 566 6,797 1,366 318 0.814 0.640 0.955 0.736 2.864
Wallhack LogisticRegression 714 7,629 1,116 253 0.859 0.738 0.968 0.805 2.861 943 5,622 3,627 82 0.639 0.920 0.986 0.764 1.420
Naive Bayes 604 7,677 1,068 363 0.853 0.625 0.955 0.751 2.741 891 6,103 3,146 134 0.681 0.869 0.979 0.765 1.545
RandomForest 967 8,745 0 0 1.000 1.000 1.000 1.000 6.713 751 8,083 1,166 274 0.860 0.733 0.967 0.803 2.872
SVM 548 8,323 422 419 0.913 0.567 0.952 0.759 4.378 817 7,227 2,022 208 0.783 0.797 0.972 0.789 2.065
MLP 472 8,489 256 495 0.923 0.488 0.945 0.729 5.186 788 7,726 1,523 237 0.829 0.769 0.970 0.802 2.471
GaussianNB 604 7,677 1,068 363 0.853 0.625 0.955 0.751 2.741 891 6,103 3,146 134 0.681 0.869 0.979 0.765 1.545
QDA 493 8,276 469 474 0.903 0.510 0.946 0.728 4.057 815 7,270 1,979 210 0.787 0.795 0.972 0.791 2.095
RevStats 604 8,241 504 363 0.911 0.625 0.958 0.783 4.143 855 6,988 2,261 170 0.763 0.834 0.976 0.795 1.946
1CT: Cheat type; 2OEI: Operational Efficiency Index.
6.2 Experimental Settings and Ground Truth
@@
The demos are open-sourced and acquired from a well-known and recognized CS:GO platform [^3] Link. We anonymize the platform’s URL and name. The initial data extraction utilizes an open-source library awpy [67] to parse the raw data in demos into JSON files, which contain comprehensive information for the following feature constructions.
Table 1 is the dataset description. Figure 5 illustrates the experiments’ process. We evaluate Hawk with two of the most prevalent cheat types, aimbot and wallhack. For aimbot, we set one collection window (16 days in total). For wallhack, we set two collection windows (27 days in total, we explain the reason for window inconsistency in Section 6.7). All demos are real-world ranking matches within the collection windows. Therefore, the cheats used by the player are black-boxes for anti-cheating. Both datasets contain different levels of cheating sophistication.
The dates of cheaters’ label acquisition both end on the last day of data collection in each set (i.e., first labeling), this label is used for training and most of the evaluations in Figure 5. The official inspections are continuous, including multiple industry-in-use proprietary closed-source anti-cheats and manual checks. The officials update the ban list daily. However, due to the long ban cycle of the official inspections, most cheaters will not get banned on the day they cheat. Therefore, after 65 days of the first labeling, we once again acquired the labels (i.e., second labeling) of the whole datasets to screen out cheaters who had not been labeled correctly in the first labeling for further analysis on the false positives (the increased cheater numbers are shown in Table 4).
@@
Although the two datasets contain different types of cheat, the official in-use anti-cheats can not label the specific subcategories, but can only annotate the broad categories of cheats. Cheating sophistication is subjective and difficult to quantify. Hence, even though we want to provide further information, it is infeasible to provide a more detailed breakdown of the datasets. Besides, due to the low recall and accuracy of the current in-use anti-cheats, there will be inevitable noise (cheaters who have evaded the official anti-cheats) in the negative labels within the ground truth.
Figure 5: Experiments’ process and labeling details.
6.3 Evaluation Metrics
Our chosen evaluation metrics are more skewed towards reflecting how well the cheaters can be identified. We use the following metrics in Appendix F and Operational Efficiency Index (OEI) to distinguish between cheating and normal behaviors. OEI is a new metric for evaluating the efficiency of industrial operations. Operational Efficiency Index (OEI), represented in Equation 10, where $N$ represents the total number of players. It reflects the operational efficiency gains in manual review processes. A higher OEI indicates a greater reduction in manual verification and a more effective system.
$OEI=\left(\frac{N}{TP+FP}\right)\cdot(Recall\cdot NPV)$ (10)
6.4 Ablation Study Results
To answer RQ.1, we conduct an ablation study on Hawk shown in Table 2. We set the Task-Specified Threshold Optimizer to achieve the highest accuracy with at least 70% recall for detecting aimbot, and the highest AUC-ROC for detecting wallhack, illustrating different anti-cheat scenarios. Overall, Hawk excels in identifying authentic cheaters, evidenced by the highest recall and NPV across all evaluations. For aimbot detection, RevStats enhances accuracy by effectively filtering normal players in both validation and test sets. In wallhack detection, RevStats excels in the validation set, while ExSPC leads in the test set. Although some components like RevStats and ExSPC outperform Hawk in accuracy, their lower recall (up to 64%) limits their effectiveness, affirming Hawk as the optimal choice. Notably, the Wallhack dataset yields better overall performance than the Aimbot dataset. We believe this is due to the latter’s diverse subcategories of aimbots. In Table 5 and Table 3, we present RevPov’s and RevStats’s model selection results among hundreds of tested models, and promising models are demonstrated. RevPov outperforms other models with balanced, consistent accuracy and recall. Random Forest outperformed other models under the sampled data in RevPov. Thus, we adopt multi-forest and multi-subsampling with Random Forest. RevStats outperforms individual classifiers in balancing recall and accuracy by capitalizing on the strengths of each classifier while mitigating their weaknesses. The training and validation loss for ExSPC in Figure 9 and Figure 10 (further discussed in Appendix A).
@@ 6.4 Ablation Study Results
Table 4: The manual average ban duration comparison and the remediation rate among Hawk’s false positives after the first labeling.
CT1 Hawk 2 Off.3 Validation Test
#MTPs4 Rmd%5 #IncrPs6 #MTPs4 Rmd%5 #IncrPs6
@@
1CT: Cheat type;
2
Hawk: The process time of Hawk varies with match length, Hawk’s detailed overheads are evaluated in Section 6.9;
3
Off.: The average ban duration of official manual inspections per match;
4
#MTPs: The number of missed true positives among the false positives;
5
Rmd%: The percentage of #MTPs in the number of false positives;
6
#IncrPs: The number of increased positives (cheaters) labeled by the officials after the second labeling.
Table 5: Comparison result on the different model selection of RevPov. N/A in OEI indicates the denominator is 0.
CT1 Model Validation Test
TP TN FP FN Accuracy Recall NPV AUC-ROC OEI2 TP TN FP FN Accuracy Recall NPV AUC-ROC OEI2
Aimbot REPTree [21] 19 7,660 37 845 0.897 0.022 0.901 0.509 3.028 33 8,118 45 851 0.901 0.037 0.905 0.516 3.919
RandomTree 123 6,914 783 741 0.822 0.142 0.903 0.520 1.215 125 7,399 764 759 0.832 0.141 0.907 0.524 1.305
RseslibKNN [26] 40 7,460 237 824 0.876 0.046 0.901 0.508 1.289 33 7,868 295 851 0.873 0.037 0.902 0.501 0.929
SPAARC [68] 0 7,697 0 864 0.899 0.000 0.899 0.500 N/A 0 8,163 0 884 0.902 0.000 0.902 0.500 N/A
RandomForest 16 7,665 32 848 0.897 0.019 0.900 0.507 2.974 33 8,128 35 851 0.902 0.037 0.905 0.517 4.496
Bagging [6] 18 7,666 31 846 0.898 0.021 0.901 0.508 3.278 30 8,129 34 854 0.902 0.034 0.905 0.515 4.341
VFI [20] 822 445 7,252 42 0.148 0.951 0.914 0.505 0.922 848 472 7,691 36 0.146 0.959 0.929 0.509 0.944
JRip [10] 21 7,663 34 843 0.898 0.024 0.901 0.510 3.408 42 8,116 47 842 0.902 0.048 0.906 0.521 4.376
MODLEM [57] 19 7,588 109 845 0.889 0.022 0.900 0.504 1.323 34 8,029 134 850 0.891 0.038 0.904 0.511 1.873
CHIRP [65] 0 7,697 0 864 0.899 0.000 0.899 0.500 N/A 0 8,163 0 884 0.902 0.000 0.902 0.500 N/A
CSForest [52] 50 7,292 405 814 0.858 0.058 0.900 0.503 0.980 58 7,755 408 826 0.864 0.066 0.904 0.508 1.151
RevPov 458 4,796 2,901 406 0.614 0.530 0.922 0.577 1.246 481 5,088 3,075 403 0.616 0.544 0.927 0.584 1.283
Wallhack REPTree 47 8,641 104 920 0.895 0.049 0.904 0.518 2.825 66 9,140 109 959 0.896 0.064 0.905 0.526 3.421
RandomTree 178 8,006 738 789 0.843 0.184 0.910 0.550 1.776 198 8,433 816 827 0.840 0.193 0.911 0.552 1.782
RseslibKNN 89 8,515 230 878 0.886 0.092 0.907 0.533 2.540 120 8,939 310 905 0.882 0.117 0.908 0.542 2.540
SPAARC 69 8,679 66 898 0.901 0.071 0.906 0.532 4.652 87 9,164 85 938 0.900 0.085 0.907 0.538 4.599
RandomForest 27 8,713 32 940 0.900 0.028 0.903 0.512 4.149 35 9,212 37 990 0.900 0.034 0.903 0.515 4.400
Bagging 26 8,705 40 941 0.899 0.027 0.902 0.511 3.571 37 9,201 48 988 0.899 0.036 0.903 0.515 3.940
VFI 950 563 8,182 17 0.156 0.982 0.971 0.523 1.014 1,008 689 8,560 17 0.166 0.983 0.976 0.529 1.031
JRip 36 8,670 75 931 0.896 0.037 0.903 0.514 2.941 71 9,179 70 954 0.900 0.069 0.906 0.531 4.572
MODLEM 35 8,687 58 932 0.898 0.036 0.903 0.515 3.414 56 9,171 78 969 0.898 0.055 0.904 0.523 3.789
CHIRP 6 8,738 7 961 0.900 0.006 0.901 0.503 4.176 8 9,242 7 1,017 0.900 0.008 0.901 0.504 4.816
CSForest 0 8,745 0 967 0.900 0.000 0.900 0.500 N/A 0 9,249 0 1,025 0.900 0.000 0.900 0.500 N/A
RevPov 573 6,081 2,664 394 0.685 0.593 0.939 0.644 1.670 627 6,364 2,885 398 0.680 0,612 0.941 0.650 1.684
1CT: Cheat type; 2OEI: Operational Efficiency Index.

### v2

@@ VI Experiments
TABLE III: Incomparability to prior works.
Work Task Data Extraction Targeting Hardware Constraint Code Availability
Detection Prevention Runtime Replay file Aimbot Wallhack
BotScreen [6] ✓ (deprecated) Intel SGX (deprecated)
BlackMirror [9] Intel SGX (deprecated)
Invisibility Cloak [8] None
Prior works [10, 11, 12, 13] ✓ (unavailable) None
Hawk None
TABLE IV: Ablation study results on each of the subsystems.
CT System Validation Test
TP TN FP FN Accuracy Recall NPV AUC-ROC FPR TP TN FP FN Accuracy Recall NPV AUC-ROC FPR
Aimbot RevPov 458 4,796 2,901 406 0.614 0.530 0.922 0.577 0.378 481 5,088 3,075 403 0.616 0.544 0.927 0.584 0.377
RevStats 453 7,290 407 411 0.904 0.524 0.947 0.736 0.053 566 6,797 1,366 318 0.814 0.640 0.955 0.736 0.167
ExSPC 553 6,187 1,510 311 0.787 0.640 0.952 0.779 0.196 558 6,616 1,547 326 0.793 0.631 0.953 0.786 0.190
RevPov+RevStats 648 4,603 3,094 216 0.613 0.750 0.955 0.760 0.402 691 4,364 3,799 193 0.559 0.782 0.958 0.750 0.465
RevPov+ExSPC 611 5,733 1,964 253 0.741 0.707 0.958 0.781 0.255 608 6,082 2,081 276 0.739 0.688 0.957 0.784 0.255
RevStats+ExSPC 605 5,828 1,869 259 0.751 0.700 0.957 0.798 0.243 639 5,877 2,286 245 0.720 0.723 0.960 0.796 0.280
Hawk 614 5,803 1,894 250 0.750 0.711 0.959 0.732 0.246 642 5,837 2,326 242 0.716 0.726 0.960 0.721 0.285
Wallhack RevPov 573 6,081 2,664 394 0.685 0.593 0.939 0.644 0.305 627 6,364 2,885 398 0.680 0.612 0.941 0.650 0.312
RevStats 604 8,241 504 363 0.911 0.625 0.958 0.783 0.058 855 6,988 2,261 170 0.763 0.834 0.976 0.795 0.245
ExSPC 760 7,258 1,487 207 0.826 0.786 0.972 0.885 0.170 501 8,484 765 524 0.875 0.489 0.942 0.840 0.083
RevPov+RevStats 604 8,241 504 363 0.911 0.625 0.958 0.798 0.058 855 6,988 2,261 170 0.763 0.834 0.976 0.816 0.244
RevPov+ExSPC 806 6,900 1,845 161 0.793 0.834 0.977 0.885 0.211 573 8,257 992 452 0.859 0.559 0.948 0.840 0.107
RevStats+ExSPC 791 7,110 1,635 176 0.814 0.818 0.976 0.889 0.187 868 6,907 2,342 157 0.757 0.847 0.978 0.859 0.253
Hawk 798 7,033 1,712 169 0.806 0.825 0.977 0.815 0.196 869 6,902 2,347 156 0.756 0.848 0.978 0.797 0.254
@@
TABLE V: Selected comparison result between different superior baseline models and RevStats.
CT Model Validation Test
TP TN FP FN Accuracy Recall NPV AUC-ROC FPR TP TN FP FN Accuracy Recall NPV AUC-ROC FPR
Aimbot LogisticRegression [46] 570 6,204 1,493 294 0.791 0.660 0.955 0.733 0.194 713 4,897 3,266 171 0.620 0.807 0.966 0.703 0.400
Naive Bayes [49] 467 6,740 957 397 0.842 0.874 0.944 0.708 0.124 625 5,928 2,235 259 0.724 0.707 0.958 0.717 0.274
RandomForest [47] 415 7,694 3 449 0.947 0.480 0.945 0.740 0.000 427 7,634 529 457 0.891 0.483 0.944 0.709 0.065
SVM [48] 309 7,615 82 555 0.926 0.358 0.932 0.673 0.011 405 7,665 498 479 0.892 0.458 0.941 0.699 0.061
MLP [45] 858 7,697 0 6 0.999 0.993 0.999 0.997 0.000 508 7,325 838 376 0.866 0.575 0.951 0.736 0.103
GaussianNB [50] 467 6,740 957 397 0.842 0.541 0.944 0.708 0.124 625 5,928 2,235 259 0.724 0.707 0.958 0.717 0.274
QDA [51] 350 7,181 516 514 0.880 0.405 0.933 0.669 0.067 534 6,770 1,393 350 0.807 0.604 0.951 0.717 0.171
RevStats 453 7,290 407 411 0.904 0.524 0.947 0.736 0.053 566 6,797 1,366 318 0.814 0.640 0.955 0.736 0.167
Wallhack LogisticRegression 714 7,629 1,116 253 0.859 0.738 0.968 0.805 0.128 943 5,622 3,627 82 0.639 0.920 0.986 0.764 0.392
Naive Bayes 604 7,677 1,068 363 0.853 0.625 0.955 0.751 0.122 891 6,103 3,146 134 0.681 0.869 0.979 0.765 0.340
RandomForest 967 8,745 0 0 1.000 1.000 1.000 1.000 0.000 751 8,083 1,166 274 0.860 0.733 0.967 0.803 0.126
SVM 548 8,323 422 419 0.913 0.567 0.952 0.759 0.048 817 7,227 2,022 208 0.783 0.797 0.972 0.789 0.219
MLP 472 8,489 256 495 0.923 0.488 0.945 0.729 0.029 788 7,726 1,523 237 0.829 0.769 0.970 0.802 0.165
GaussianNB 604 7,677 1,068 363 0.853 0.625 0.955 0.751 0.122 891 6,103 3,146 134 0.681 0.869 0.979 0.765 0.340
QDA 493 8,276 469 474 0.903 0.510 0.946 0.728 0.054 815 7,270 1,979 210 0.787 0.795 0.972 0.791 0.214
RevStats 604 8,241 504 363 0.911 0.625 0.958 0.783 0.058 855 6,988 2,261 170 0.763 0.834 0.976 0.795 0.245
VI-A Incomparability to Prior Works
First, most prior works are not open-sourced as listed in Table I. Second, the data collection method is incompatible. Hawk relies on post-game replay files, whereas prior studies modified the game engine or applied mods during gameplay. Prior works are unable to extract data from the replay files. Third, installation on commercial clients or servers is infeasible. Only if the prior designs are implemented before the data collection, the data can then be gathered and evaluated. However, the partner platform prohibits installing unverified programs on clients or servers due to legal restrictions. In addition to the aforementioned reasons with commonality, we also further state below the specific reasons for the detailed incomparability between Hawk and the SOTA designs. There are three aspects of incomparability with BotScreen [6]. First, BotScreen’s data extraction is highly coupled with the game version and deprecated. The authors of BotScreen confirmed through email that their data extraction does not apply to the current game versions. Second, the deprecation of Intel SGX [14, 15] prevents us from conducting comparative experiments. Third, it can only detect pure aimbot, making the comparison less significant. Invisibility Cloak [8] and BlackMirror [9] are prevention designs only for counteracting computer-vision-based aimbots and wallhacks, respectively. Both studies are only functional before the cheats happen.
Figure 6: Recall and FPR performance comparison among Hawk-Recall priority mode, Hawk-Balance mode, and our partner in-use anti-cheat [^4] Commercial anti-cheat (detailed in Section VIII), not open-source, leverages reverse engineering on the known cheats to detect cheating signatures.. Hawk’s different modes are introduced and evaluated in Section V-D and Section VI-F.
Whereas Hawk and other prior works are detection systems that function after cheating. We detailed the incomparability to prior works in Table III. We manage to compare Hawk with our partner’s industrial in-use anti-cheat in footnote 4 and reproduce two server-side works [11, 12] for further comparison in Table IX.
VI-B Experimental Settings and Ground Truth
@@
The demos are open-sourced and acquired from a well-known CS:GO platform5555EPlay: https://www.5eplay.com/.. The initial data extraction utilizes a public library awpy [52] to parse the demo into a JSON file, which contains the raw data for Hawk’s feature constructions.
Table II is the dataset description. Figure 5 illustrates our experimental process. We evaluate Hawk with two of the most prevalent cheat types, aimbot and wallhack. For aimbot, data is collected over a single 16-day window, while for wallhack, two collection windows totaling 27 days are used (the rationale for this inconsistency is explained in Section VI-G). All demos are from real-world ranked matches within these windows, ensuring that the cheats used by players remain black-box to the anti-cheat system. Both datasets naturally contain varying levels of cheating sophistication. The cheaters’ labels in the first labeling are acquired on the last day of data collection for each set. These labels are used for training and most evaluations, as illustrated in Figure 5. Official inspections, which include proprietary closed-source anti-cheat systems and manual checks, are conducted continuously, with daily updates to the ban list. However, due to the lengthy ban cycle, most cheaters are not banned immediately after cheating. To address this, a second labeling is performed 65 days after the first labeling to identify cheaters who were initially mislabeled. This allowed us to refine our analysis of false positives, with the updated cheater counts detailed in Table VI.
Official inspections employ a combination of proprietary anti-cheat software, developed by 5Eplay, and manual verification. The anti-cheat system operates by installing a kernel-level driver to scan system memory and active processes for known cheat signatures. A significant drawback of this method is its inherent reliance on a blacklist, making it one step behind the cheating developers. Modern cheats frequently update their code and employ advanced obfuscation techniques to evade signature-based detection. When a potential violation is identified, the system forwards the detection results to human reviewers for final determination. Note that despite 5EPlay tried their best to annotate the samples as accurately as possible. This process inherently and inevitably introduces a small amount of false negatives (i.e., the cheaters who escape the inspections) after the second labeling due to the current limitation of anti-cheat low performance. This is an unsolvable dilemma for this area. All positive samples (cheaters) are accurately labeled.
The increase of cheaters between the first and second labeling is indicated in Table VI’s #IncrPs. Based on 5EPlay’s statistics and the case study in Section VI-H, there are approximately 1.5%-2% and 3%-4% uncaught cheaters in the negative samples after the second labeling across different sets in the Aimbot and Wallhack datasets, respectively.
@@
Although the two datasets include different cheat types, the official in-use anti-cheat systems cannot label specific subcategories; they only annotate broad cheat categories. Quantifying cheating sophistication is inherently subjective and challenging, making it infeasible to provide a more detailed breakdown of the datasets. Additionally, due to the poor performance of current anti-cheat systems, the ground truth inevitably contains noise (cheaters who evade detection) within the negative labels. But to the best of our knowledge, this is the most realistic and well-annotated dataset.
To evaluate how effectively cheaters are identified, we use the metrics, accuracy, recall, NPV, AUC-ROC, and FPR, defined in Supplemental Material G.
@@
TABLE VI: The manual average ban duration comparison and the remediation rate among Hawk’s false positives after the first labeling.
CT Hawk1 Off.2 Validation Test
#MTPs3 Rmd%4 #IncrPs5 #MTPs Rmd% #IncrPs
@@
1Hawk: The process time of Hawk varies with match length, Hawk’s detailed overheads are evaluated in Section VI-I; 2Off.: The average ban duration of official manual inspections per match; 3#MTPs: The number of missed true positives among the false positives; 4Rmd%: The percentage of #MTPs in the number of false positives; 5#IncrPs: The number of increased positives (cheaters) labeled by the officials after the second labeling.
TABLE VII: Comparison result on the different baseline models of RevPov.
CT Model Validation Test
TP TN FP FN Accuracy Recall NPV AUC-ROC FPR TP TN FP FN Accuracy Recall NPV AUC-ROC FPR
Aimbot REPTree [53] 19 7,660 37 845 0.897 0.022 0.901 0.509 0.005 33 8,118 45 851 0.901 0.037 0.905 0.516 0.006
RandomTree 123 6,914 783 741 0.822 0.142 0.903 0.520 0.102 125 7,399 764 759 0.832 0.141 0.907 0.524 0.094
RseslibKNN [54] 40 7,460 237 824 0.876 0.046 0.901 0.508 0.031 33 7,868 295 851 0.873 0.037 0.902 0.501 0.036
SPAARC [55] 0 7,697 0 864 0.899 0.000 0.899 0.500 0.000 0 8,163 0 884 0.902 0.000 0.902 0.500 0.000
RandomForest [47] 16 7,665 32 848 0.897 0.019 0.900 0.507 0.004 33 8,128 35 851 0.902 0.037 0.905 0.517 0.004
Bagging [56] 18 7,666 31 846 0.898 0.021 0.901 0.508 0.004 30 8,129 34 854 0.902 0.034 0.905 0.515 0.004
VFI [57] 822 445 7,252 42 0.148 0.951 0.914 0.505 0.942 848 472 7,691 36 0.146 0.959 0.929 0.509 0.942
JRip [58] 21 7,663 34 843 0.898 0.024 0.901 0.510 0.004 42 8,116 47 842 0.902 0.048 0.906 0.521 0.006
MODLEM [59] 19 7,588 109 845 0.889 0.022 0.900 0.504 0.014 34 8,029 134 850 0.891 0.038 0.904 0.511 0.016
CSForest [60] 50 7,292 405 814 0.858 0.058 0.900 0.503 0.053 58 7,755 408 826 0.864 0.066 0.904 0.508 0.050
RevPov 458 4,796 2,901 406 0.614 0.530 0.922 0.577 0.377 481 5,088 3,075 403 0.616 0.544 0.927 0.584 0.377
Wallhack REPTree 47 8,641 104 920 0.895 0.049 0.904 0.518 0.012 66 9,140 109 959 0.896 0.064 0.905 0.526 0.012
RandomTree 178 8,006 738 789 0.843 0.184 0.910 0.550 0.084 198 8,433 816 827 0.840 0.193 0.911 0.552 0.088
RseslibKNN 89 8,515 230 878 0.886 0.092 0.907 0.533 0.026 120 8,939 310 905 0.882 0.117 0.908 0.542 0.034
SPAARC 69 8,679 66 898 0.901 0.071 0.906 0.532 0.008 87 9,164 85 938 0.900 0.085 0.907 0.538 0.009
RandomForest 27 8,713 32 940 0.900 0.028 0.903 0.512 0.004 35 9,212 37 990 0.900 0.034 0.903 0.515 0.004
Bagging 26 8,705 40 941 0.899 0.027 0.902 0.511 0.005 37 9,201 48 988 0.899 0.036 0.903 0.515 0.005
VFI 950 563 8,182 17 0.156 0.982 0.971 0.523 0.936 1,008 689 8,560 17 0.166 0.983 0.976 0.529 0.925
JRip 36 8,670 75 931 0.896 0.037 0.903 0.514 0.009 71 9,179 70 954 0.900 0.069 0.906 0.531 0.008
MODLEM 35 8,687 58 932 0.898 0.036 0.903 0.515 0.007 56 9,171 78 969 0.898 0.055 0.904 0.523 0.008
CSForest 0 8,745 0 967 0.900 0.000 0.900 0.500 0.000 0 9,249 0 1,025 0.900 0.000 0.900 0.500 0.000
RevPov 573 6,081 2,664 394 0.685 0.593 0.939 0.644 0.305 627 6,364 2,885 398 0.680 0,612 0.941 0.650 0.312